The profit on a watch is given by P = x2 – 13x – 80, where x is the number of watches sold per day. How many watches were sold on a day when there was a $50 loss?
Found 2 solutions by checkley75, ankor@dixie-net.com:
Answer by checkley75(3666)

-50=x^2-13x-80
x^2-13x-80+50
x^2-13x-30=0
(x-15)(x+2)=0
x-15=0
x=15 answer.
x+2=0
x=-2 can't sell -2 watches. not an answer.

Answer by ankor@dixie-net.com(12704)

The profit on a watch is given by P = x^Ch2 – 13x – 80, where x is the number of watches sold per day. How many watches were sold on a day when there was a $50 loss?
:
x^2 - 13x - 80 = -50; negative because of a loss of 50
:
x^2 - 13x - 80 + 50 = 0; add 50 to both sides
:
x^2 - 13x - 30 = 0
:
Easily factors to:
(x-15) (x+2) = 0
:
x = -2
and
x = +15 watches sold for a loss of $50
:
:
Check; substitute 15 for x in the original equation:
15^2 - 13(15) - 80 =
225 - 195 - 80 = -50